Editor of school newspaper must reduce the size of a graph to fit one column.
Original is 2" x 2", and the scale factor from the original to the reduced graph is 8:3. Find the dimensions of the graph as it will appear in one column of newspaper.
I have tried: 2/8 2/3 = 1/4 . 2/3 = 3/8 ???? Please help.

Answer by stanbon(75887)

1st Method:
If you are supposed to be reducing the dimensions you get:
Proportion:
8/3 = 2/x
Invert to get:
x/2 = 3/8
Multiply both sides by 2 to get:
x = 3/4"
Ans: 3/4" by 3/4"
======================
2nd Method: But if you are supposed to be reducing the area, proceed
as follows:
-------
Original area = 2*2 = 4 sq. in.
----
Use proportion:
x/4 = 3/8
x = 3/2 sq. in
---
Then the dimensions would be sqrt(3/2) by sqrt(3/2)
or (1/2)sqrt(6) by (1/2)sqrt(6)
===================================
Take your choice.
-----------------------------------
Cheers,
Stan H.
Cheers--,
Stan H.